# Janus: Resilient and Adaptive Data Transmission for Enabling Timely and Efficient Cross-Facility Scientific Workflows

### v1

@@ 5.2.3. Data Transfer under Static Packet Loss Rates
Figure 2. Total time for transferring data with guaranteed error bound under static packet loss rates
While real-world network packet loss rates fluctuate over time, we conducted simulation-based evaluations under static packet loss rates to validate the correctness of the models we developed in Section 3.
First, we evaluated the data transfer process for ensuring an error-bound guarantee on the received data. Assuming the user specifies that the error in the received data must remain below $0.00001$, all four levels of the refactored NYX cosmology simulation data must be successfully received to meet this requirement. We simulated the transfer of these four levels using the TCP protocol. We also simulated an alternative UDP-based protocol that incorporates erasure coding with a static number of parity fragments for each FTG, denoted as $m$, alongside passive retransmission to ensure fault tolerance. For each possible value of $m$, we conducted simulations, recording the total time required to complete the transfer of all four levels. Furthermore, we computed the expected total transfer time, $E[T_{total}]$, using the equations presented in Section 3 and compared the calculated results with the simulation outcomes. The evaluation results are presented in Fig. 2, providing the following insights: 1) When the TCP protocol is used, the data transmission time increases significantly as the packet loss rate rises. 2) For the UDP-based protocol that integrates erasure coding and passive retransmission, the total transmission time derived from the theoretical equations in Section 3 aligns closely with the simulation results. This alignment confirms that our model effectively captures the impact of packet loss, parity fragment generation, and retransmission on overall transmission performance. 3) When the packet loss rate is low (Fig. 2(a)), increasing the number of parity fragments per FTG negatively affects performance due to the added overhead. However, as the packet loss rate increases (Fig. 2(b) and Fig. 2(c)), there exists an optimal value of $m$ that minimizes the total transmission time.
@@
Figure 3. Error bounds of data received within guaranteed transmission time under static packet loss rates
Second, we evaluated the data transfer process under a guaranteed transmission time constraint, which enables users to receive data within a specified time frame while accepting the trade-off that the error bound of the received data may not be strictly guaranteed. As shown in Fig. 2, the minimum transmission times required to transfer all levels of the refactored data under different packet loss rates were 378.03 seconds for $\lambda=19$, 401.11 seconds for $\lambda=383$, and 429.75 seconds for $\lambda=957$. These values were subsequently used as the transmission time constraint $\tau$ in solving the optimization model presented in Equation 12. The optimization results indicate that to minimize the error bound of the received data, the optimal number of parity fragments per FTG at all four levels should be set as follows: for $\lambda=19$, the values are $m_{1}=5$, $m_{2}=4$, $m_{3}=2$, and $m_{4}=0$; for $\lambda=383$, the values are $m_{1}=8$, $m_{2}=7$, $m_{3}=7$, and $m_{4}=0$; and for $\lambda=957$, the values are $m_{1}=12$, $m_{2}=11$, $m_{3}=11$, and $m_{4}=0$. To evaluate the effectiveness of these fault-tolerance configurations, we conducted simulations using the derived configurations, running each scenario 100 times to account for the nondeterministic nature of data errors. We recorded the number of instances where the received data satisfied different error bounds, such as $\varepsilon_{0}$, $\varepsilon_{1}$, and so on. This approach allowed us to quantify the frequency of achieving varying levels of data accuracy under the given configurations. Additionally, to highlight the advantages of the optimized configurations, we performed simulations with alternative configurations where the number of parity fragments per FTG was set uniformly across all four levels. As illustrated in Fig. 3, the optimized configurations consistently ensured that the transmission time remained within the specified constraint while achieving a high likelihood of maintaining a low error bound ($\varepsilon_{3}$) in the received data. In contrast, the alternative configurations either met the time constraint but failed to recover the data entirely (resulting in an error bound of $\varepsilon_{0}$), or exceeded the specified transmission time constraint.
5.2.4. Data Transfer under Time-Varying Packet Loss Rates
When the packet loss rate changes over time, a static fault-tolerance configuration fails to achieve optimal transmission performance. We conducted simulation-based evaluations to assess the effectiveness of the proposed data transfer protocols in dynamically adjusting fault-tolerance configurations in response to time-varying packet loss rates.
First, we conducted simulations to evaluate the performance of three data transfer approaches under a guaranteed error bound: the TCP protocol, a UDP-based protocol incorporating erasure coding with a static number of parity fragments per FTG, and the proposed adaptive protocol shown in algorithm 1. As illustrated in Fig. 4, the adaptive protocol consistently outperforms the other two approaches under time-varying packet loss rates. Notably, the adaptive protocol completes the data transmission in 388.8 seconds, reducing the transmission time by 30 seconds compared to the minimum time achieved by the static fault-tolerance configuration.
Figure 4. Total time for transferring data with guaranteed error bound under time-varying packet loss rates
Second, we conducted simulations to evaluate the proposed protocol in algorithm 2 for transferring data within a guaranteed time constraint under time-varying packet loss rates. Since the minimum transmission time for all data levels using the protocol in algorithm 1 under time-varying packet loss rates is 388.8 seconds, we selected this duration as the time constraint. This choice reflects a scenario where the user is willing to trade some accuracy of the received data for potentially reduced transmission time. To determine the optimal fault-tolerance configuration under this constraint, we solved the optimization model in Equation 12 using static packet loss rates. The results indicate that to meet the time constraint while minimizing the error bound, the optimal number of parity fragments per FTG at all four levels should be configured as follows: for $\lambda=19$, the values are $m_{1}=9$, $m_{2}=6$, $m_{3}=4$, and $m_{4}=0$; for $\lambda=383$, the values are $m_{1}=16$, $m_{2}=8$, $m_{3}=0$, and $m_{4}=0$; and for $\lambda=957$, the values are $m_{1}=15$, $m_{2}=9$, $m_{3}=0$, and $m_{4}=0$. We performed 100 simulation runs using these static fault-tolerance configurations under time-varying packet loss rates. Additionally, we ran another 100 simulations using the adaptive protocol in algorithm 2 under the same conditions, recording the number of instances where the received data met various error bounds. As shown in Fig. 5, both the static and adaptive fault-tolerance configurations successfully met the time constraint. However, the adaptive configuration consistently resulted in received data satisfying lower error bounds more frequently than the static configurations, demonstrating its effectiveness in improving data accuracy while maintaining the required transmission time under fluctuating packet loss rates.
Figure 5. Error bounds of data received within guaranteed transmission time under time-varying packet loss rates
5.3. Real Network Testing
5.3.1. C++ Implementation of Adaptive Data Transfer Protocols
@@
To evaluate the performance of data transmission over real wide-area networks, we developed a prototype of the proposed adaptive data transfer protocols, as described in algorithm 1 and algorithm 2, using C++. Our implementation leverages the Boost.Asio (Boost.Asio, 2025) library for efficient UDP-based data transmission. Additionally, we utilize the Protocol Buffers (Protobuf) library (Protocol Buffers, 2025) to structure packet contents, embedding essential erasure coding metadata—such as the level and FTG each fragment belongs to, as well as its corresponding redundancy level ($m$)—alongside the actual data or parity fragment. The prototype, comprising over 1,600 lines of C++ code, provides a scalable and flexible foundation for evaluating the proposed adaptive transfer mechanisms in real-network environments.
5.3.2. Testing Results
@@
Figure 6. Total time for transferring data with guaranteed error bound using different data transfer protocols
In each test run, we conducted several measurements. First, we used the native TCP protocol to transfer all levels of the refactored Nyx data and recorded the total transmission time. Next, we employed the Globus (Globus, 2025) data transfer service to transfer the same data and recorded the transmission time. Then we used the adaptive protocol described in algorithm 1 to transmit the data, ensuring an error bound guarantee of $\varepsilon_{4}$ and requiring the successful receipt of all levels. Finally, we applied the adaptive protocol in algorithm 2 to transfer the data within a fixed time constraint, set to 90% of the transmission time achieved by the protocol in algorithm 1. To account for variations in packet loss rates, we conducted five test runs at different times and on different days.
Test ID Time Constraint(s) Error Bound
1 367.06 $\varepsilon_{2}$
2 392.08 $\varepsilon_{2}$
3 339.32 $\varepsilon_{2}$
4 371.08 $\varepsilon_{2}$
5 348.66 $\varepsilon_{1}$
@@
Table 2. Error bounds of data received within guaranteed transmission time
As shown in Fig. 6, when all levels of the Nyx data must be received, the transmission time for native TCP and Globus varies significantly across the five test runs, emphasizing their sensitivity to packet loss. In contrast, our adaptive data transfer protocol not only reduces transmission time substantially but also offers greater stability. Furthermore, as demonstrated in Table 2, when using the adaptive protocol described in algorithm 2 to meet the selected time constraints, four out of five test runs achieved data within the error bound $\varepsilon_{2}$, while one run resulted in data within the error bound $\varepsilon_{1}$. This shows that by utilizing algorithm 2, we can guarantee a shorter transmission time, albeit with a slight sacrifice in the accuracy of the received data.
6. Related Work
@@
UDP, a connectionless and low-latency transport protocol, lacks built-in error correction mechanisms (Postel, 1980), making it particularly susceptible to packet loss in high-performance and wide-area network environments. To address this limitation, many studies have explored lightweight and adaptive flow control mechanisms to improve the reliability of UDP-based data transfer. One notable approach is Performance-Adaptive UDP (PA-UDP), a high-speed data transfer protocol proposed by (Eckart et al., 2008). PA-UDP enables dynamic flow control based on the sending system’s performance characteristics, allowing UDP packets to be sent asynchronously without blocking TCP transmission. The SDUDP framework (Wang et al., 2017) takes advantage of Software-Defined Networking (SDN)’s capabilities, such as centralized flow control and real-time packet monitoring, to achieve reliable UDP data transfer. Beyond SDUDP, various studies have explored Reliable UDP (RUDP) protocols designed to enhance UDP’s performance while maintaining its low-latency advantage over TCP. These protocols typically introduce mechanisms such as selective acknowledgments, forward error correction (FEC), and retransmission strategies to mitigate packet loss. However, many RUDP implementations require modifications to end-host systems, making deployment challenging in large-scale networks (Tran and Choi, 2007; Masirap et al., 2016; Wang et al., 2006). In addition, QUIC (Quick UDP Internet Connections), a transport layer protocol developed by Google, uses mechanisms like congestion control and loss recovery to enhance packet delivery (Langley et al., 2017).
@@
To further optimize UDP-based data transfer, research has focused on detecting and predicting packet loss more efficiently. For instance, Wu et al. (Wu et al., 2022) proposed a novel framework, LossDetection, for real-time packet loss detection in both TCP and UDP traffic. LossDetection employs Feature-Sketch, a technique that extracts key features from packet streams for efficient and scalable loss identification. By detecting packet loss patterns dynamically, this method enables more effective retransmission strategies. Another approach to mitigating packet loss involves predictive modeling. Giannakou et al. (Giannakou et al., 2020) developed a lightweight Random Forest Regression (RFR) model to predict retransmissions in scientific data transfers. Their model builds upon existing loss detection techniques by forecasting packet loss trends and proactively adjusting transmission parameters to minimize unnecessary retransmissions.
@@
Beyond optimizing flow control mechanisms, another type of approach to improving UDP reliability involves introducing redundancy into transmitted data to tolerate packet losses. A fundamental technique used in these approaches is erasure coding, which has been extensively studied and implemented to enhance fault tolerance in both data storage and transmission. Foundational work by Reed and Solomon introduced Reed-Solomon codes (Wicker and Bhargava, 1999), which remain a gold standard for error correction. Several studies have explored the integration of erasure coding with UDP-based data transfer to improve reliability. For instance, Byers et al. (Byers et al., 1998) demonstrated that incorporating Reed-Solomon codes into UDP streaming significantly enhances packet recovery, mitigating losses common in lossy network conditions. Further research has investigated the use of fountain codes, such as Raptor codes, over UDP to facilitate reliable high-bandwidth data transfers. Mtibaa and Bestavros proposed RC-UDP (Mtibaa et al., 2018), a framework that leverages Raptor coding over UDP to achieve efficient and reliable data transmission. Their findings indicate that this approach effectively addresses packet loss without incurring significant overhead.

### v2

@@ V-B3 Data Transfer with Guaranteed Error Bound
(d) Time-varying packet loss rate
Figure 2: Total time for transferring data with guaranteed error bound under different packet loss rates
In this evaluation, we focus on the data transfer process where an error-bound must be guaranteed for the received data. Assuming the user specifies that the error in the received data must remain below $0.00001$, all four levels of the refactored Nyx cosmology simulation data must be successfully received to meet this requirement.
While packet loss rates in real-world wide-area networks vary dynamically over time, we first conducted simulation-based evaluations under static packet loss conditions to validate the analytical models developed in Section III. In these baseline experiments, we simulated the data transfer of four hierarchical levels using both the TCP protocol and a UDP-based fault-tolerant protocol. The UDP-based approach enhances traditional UDP transmission with erasure coding and passive retransmission. In this configuration, each fault-tolerance group (FTG) includes a fixed number of parity fragments, denoted as $m$, which are used to recover lost packets without requiring end-to-end acknowledgment. For each possible value of $m$, we performed multiple simulations and recorded the total time required to complete the transfer across all levels. In parallel, we computed the expected total transfer time, $E[T_{total}]$, using the analytical expressions derived in Section III, allowing for direct comparison between simulated and theoretical results.
The evaluation results are presented in Fig. 2(a)– Fig. 2(c), providing the following insights: 1) When the TCP protocol is used, the data transmission time increases significantly as the packet loss rate rises. 2) For the UDP-based protocol that integrates erasure coding and passive retransmission, the total transmission time derived from the theoretical equations in Section III aligns closely with the simulation results. This alignment confirms that our model effectively captures the impact of packet loss, parity fragment generation, and retransmission on overall transmission performance. 3) When the packet loss rate is low (Fig. 2(a)), increasing the number of parity fragments per FTG negatively affects performance due to the added overhead. However, as the packet loss rate increases (Fig. 2(b) and Fig. 2(c)), there exists an optimal value of $m$ that minimizes the total transmission time.
When the packet loss rate changes over time, a static fault-tolerance configuration often fails to achieve optimal transmission performance. To demonstrate this, we conducted simulation-based evaluations to assess the effectiveness of the proposed data transfer protocols in dynamically adjusting fault-tolerance configurations in response to time-varying packet loss rates. We compare the proposed protocols with two other baseline protocols: the TCP protocol, a UDP-based protocol incorporating erasure coding with a static number of parity fragments per FTG. As illustrated in Fig. 2(d), the adaptive protocol consistently outperforms the other two approaches under time-varying packet loss rates. Notably, the adaptive protocol completes the data transmission in 252.4 seconds, reducing the transmission time by about 20 seconds compared to the minimum time achieved by the static fault-tolerance configuration.
V-B4 Data Transfer with Guaranteed Time
@@
Figure 3: Error bounds of data received within guaranteed transmission time under different packet loss rates
In this evaluation, we focus on the data transfer process under a guaranteed transmission time constraint, which allows users to receive data within a specified time frame while accepting that the error bound of the received data may not be strictly guaranteed. This scenario reflects practical use cases where timely data delivery is prioritized over perfect accuracy.
We first conducted simulations under static packet loss rates to analyze the system’s behavior and validate the optimization model introduced in Equation 11. As shown inFig. 2(a)– Fig. 2(c), the minimum transmission times required to complete the transfer of all refactored data levels were 247.2 seconds under low packet loss ($\lambda=20$), 260.6 seconds under medium loss ($\lambda=400$), and 276.3 seconds under high loss ($\lambda=1000$). Since users in this mode are willing to trade data accuracy for faster completion, we defined transmission time constraints $\tau$ of 245 seconds, 250 seconds, and 275 seconds for the low, medium, and high loss scenarios, respectively. The optimization results indicate that minimizing the reconstruction error within these time limits requires adjusting the number of parity fragments per FTG across levels. Specifically, the optimal configurations were $m_{1}=3$, $m_{2}=1$, $m_{3}=1$, and $m_{4}=0$ for $\lambda=20$; $m_{1}=7$, $m_{2}=6$, $m_{3}=6$, and $m_{4}=0$ for $\lambda=400$; and $m_{1}=11$, $m_{2}=10$, $m_{3}=8$, and $m_{4}=0$ for $\lambda=1000$.
To evaluate the effectiveness of these configurations, we performed 100 independent simulation runs per scenario to account for the stochastic nature of packet loss and data corruption. In each run, we recorded the final error bounds of the received data, such as $\varepsilon_{0}$, $\varepsilon_{1}$, and $\varepsilon_{3}$. Aggregating results from 100 runs allows us to quantify the likelihood of achieving different levels of reconstruction accuracy under the specified time constraints. To assess the benefit of the optimized settings, we compared them with uniform configurations, where the number of parity fragments per FTG was fixed across all four levels. As shown in Fig. 3(a)– Fig. 3(c), the optimized configurations consistently met the transmission time constraint $\tau$ while maintaining a high probability of low error bounds (e.g., $\varepsilon_{3}$). In contrast, uniform configurations either satisfied the time constraint but resulted in high error bounds such as $\varepsilon_{0}$, or exceeded the transmission time limit. Note that in each run, simulations continued even if the transmission time exceeded the threshold. With a strict cut-off, transfers using uniform configurations would have a much higher likelihood of receiving data with higher error bounds than shown in Fig. 3(a)– Fig. 3(c).
Next, we conducted simulations to evaluate the proposed adaptive protocol in algorithm 2, which transfers data within a guaranteed time constraint under time-varying packet loss rates. Based on prior experiments, the minimum transmission time required to transfer all data levels using the protocol in algorithm 1 under the same conditions was 252.4 seconds. Accordingly, we set the transmission time constraint to 250 seconds to reflect the scenario in which the user is willing to trade a small degree of accuracy in the received data for faster delivery. Because the packet loss rate changes over time, no static fault-tolerance configuration can simultaneously minimize reconstruction error and satisfy the time constraint across all conditions. To provide a baseline for comparison, we also conducted simulations using the optimal static configurations previously derived for constant-loss scenarios (Fig. 3(a)Fig. 3(c)) under the same time-varying loss patterns.
We performed 100 simulation runs using these static fault-tolerance configurations and another 100 runs using the adaptive protocol described in algorithm 2. For each run, we recorded the number of instances in which the received data satisfied different error bounds. As shown in Fig. 3(d), the adaptive protocol consistently maintained transmission completion within the specified time constraint while achieving a high probability of low reconstruction error. In contrast, the static configurations either met the time constraint but yielded data with higher error bounds or failed to complete within the required time.
V-C Real Network Testing
V-C1 C++ Implementation of Adaptive Transfer Protocols
@@
To evaluate the performance of data transmission over real wide-area networks, we developed a prototype of the proposed adaptive data transfer protocols, as described in algorithm 1 and algorithm 2, using C++. Our implementation leverages the Boost.Asio [Boost.Asio] library for efficient UDP-based data transmission. Additionally, we utilize the Protocol Buffers (Protobuf) library [Protobuf] to structure packet contents, embedding essential erasure coding metadata—such as the level and FTG each fragment belongs to, as well as its corresponding redundancy level (mm)—alongside the actual data or parity fragment. The prototype, comprising over 1,600 lines of C++ code, provides a scalable and flexible foundation for evaluating the proposed adaptive transfer mechanisms in real-network environments.
V-C2 Experimental Setup
@@
Figure 4: Total time for transferring data with guaranteed error bound using different data transfer protocols
@@
Figure 5: Error bounds of data received within guaranteed transmission time
In each test run, we performed a series of measurements. First, we used the native TCP protocol to transfer all levels of the refactored Nyx dataset from the sender to the receiver and recorded the total transmission time. Next, we employed the Globus service [Globus] to transfer the same data and measured its transmission time. We then adopted the QUIC protocol [10.1145/3098822.3098842], the state-of-the-art UDP-based transport protocol with built-in fault-tolerance mechanisms for handling packet loss. Afterward, we applied the adaptive protocol described in algorithm 1 to transmit the data while ensuring an error bound guarantee of $\varepsilon_{4}$ and requiring successful receipt of all data levels. Finally, we used the adaptive protocol in algorithm 2 to transfer the data within a fixed time constraint set to 90% of the transmission time achieved by the error-bound-guaranteed protocol. To account for variability in packet loss rates, all experiments were repeated 80 times at different times of day and on different days.
V-C3 Testing Results
As shown in Fig. 4, when all levels of the Nyx data are required to be received, the transmission times of the native TCP, Globus, and QUIC protocols vary considerably across test runs, highlighting their sensitivity to packet loss. In contrast, our adaptive data transfer protocol not only achieves substantially lower transmission times but also exhibits much greater stability across runs. Moreover, as presented in Fig. 5, when employing the adaptive protocol described in algorithm 2 to satisfy the specified time constraints, all 100 test runs successfully delivered data within the error bound $\varepsilon_{3}$. In comparison, when the other protocols are used, the received data is more likely to exceed this error bound when the same time limit is reached.
VI Related Work
@@
UDP, a connectionless and low-latency transport protocol, lacks built-in error correction mechanisms [postel1980rfc], making it particularly susceptible to packet loss in high-performance and wide-area network environments. To address this limitation, many studies have explored lightweight and adaptive flow control mechanisms to improve the reliability of UDP-based data transfer. One notable approach is Performance-Adaptive UDP (PA-UDP), a high-speed data transfer protocol proposed by [eckart2008performance]. PA-UDP enables dynamic flow control based on the sending system’s performance characteristics, allowing UDP packets to be sent asynchronously without blocking TCP transmission. The SDUDP framework [wang2017sdudp] takes advantage of Software-Defined Networking (SDN)’s capabilities, such as centralized flow control and real-time packet monitoring, to achieve reliable UDP data transfer. Beyond SDUDP, various studies have explored Reliable UDP (RUDP) protocols designed to enhance UDP’s performance while maintaining its low-latency advantage over TCP. These protocols typically introduce mechanisms such as selective acknowledgments, forward error correction (FEC), and retransmission strategies to mitigate packet loss. However, many RUDP implementations require modifications to end-host systems, making deployment challenging in large-scale networks [tran2007reliable, masirap2016evaluation, wang2006reliable]. In addition, QUIC (Quick UDP Internet Connections), a transport layer protocol developed by Google, uses mechanisms like congestion control and loss recovery to enhance packet delivery [10.1145/3098822.3098842].
@@
To further optimize UDP-based data transfer, research has focused on detecting and predicting packet loss more efficiently. For instance, Wu et al. [wu2022lossdetection] proposed a novel framework, LossDetection, for real-time packet loss detection in both TCP and UDP traffic. LossDetection employs Feature-Sketch, a technique that extracts key features from packet streams for efficient and scalable loss identification. By detecting packet loss patterns dynamically, this method enables more effective retransmission strategies. Another approach to mitigating packet loss involves predictive modeling. Giannakou et al. [giannakou2020machine] developed a lightweight Random Forest Regression (RFR) model to predict retransmissions in scientific data transfers. Their model builds upon existing loss detection techniques by forecasting packet loss trends and proactively adjusting transmission parameters to minimize unnecessary retransmissions.
@@
Beyond optimizing flow control mechanisms, another type of approach to improving UDP reliability involves introducing redundancy into transmitted data to tolerate packet losses. A fundamental technique used in these approaches is erasure coding, which has been extensively studied and implemented to enhance fault tolerance in both data storage and transmission. Foundational work by Reed and Solomon introduced Reed-Solomon codes [wicker1999reed], which remain a gold standard for error correction. Several studies have explored the integration of erasure coding with UDP-based data transfer to improve reliability. For instance, Byers et al. [byers1998digital] demonstrated that incorporating Reed-Solomon codes into UDP streaming significantly enhances packet recovery, mitigating losses common in lossy network conditions. Further research has investigated the use of fountain codes, such as Raptor codes, over UDP to facilitate reliable high-bandwidth data transfers. Mtibaa and Bestavros proposed RC-UDP [mtibaa2018rc], a framework that leverages Raptor coding over UDP to achieve efficient and reliable data transmission. Their findings indicate that this approach effectively addresses packet loss without incurring significant overhead.
@@ VII Conclusion
Acknowledgment
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-23-2-0224. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.